OpenStudy (anonymous):

express all functions in terms of sin x and cos x combine the expression into a single fraction in lowest terms. 1. sec x cot x

OpenStudy (anonymous):

\[\sec(x)*\cot(x) = (1/\cos(x))*(1/\tan(x))\] Now you must remember that tan(x) = sin(x)/cos(x) and continue simplifying.

OpenStudy (anonymous):

yes i remember it

OpenStudy (anonymous):

how about this

OpenStudy (anonymous):

sin x / csc

OpenStudy (anonymous):

The answer is 1/sin(x). Work from what I gave you, and replace tan(x) with sin(x)/cos(x)...then simplify.

OpenStudy (anonymous):

about this sir COS θ = -12/13 (3RD QUADRANT) my answer is sin 13/13? is it correct?

OpenStudy (anonymous):

stephanie, please limit your questions to one per topic, this is getting very confusing. And no, that is not correct. Honestly, I suggest you first of all go read on trigonometry, because you are clearly in need of studying this subject. I can help you, but I can't be here solving your entire homework. I just feel your understanding of any of this is very flimsy...

OpenStudy (anonymous):

sorry im just asking..

OpenStudy (anonymous):

I'm ont blaming you, I am just suggesting that you should really review Trigonometry in general before attempting to do this homework. You should review the trigonometric identities, the trigonometric circle and the definitions of the each of the trigonometric functions... otherwise you won't get anywhere. Just my opinion.

OpenStudy (anonymous):

sometimes when i read it the question is not on the net or a book or sometimes the question like problem word has a trick that i cant explain.sorry dont worry this is my last Homework because next week it final exam.

OpenStudy (anonymous):

Well I imagine you have to know these things for your final exam, so I really recommend you study the subject further :P. Most of trigonometry problems like this one involve you knowing how to play around with the identities and definitions, so you really just have to know those well, then try to replace them in things and simplify.